OpenStudy (anonymous):

Using the equation you found for “7 less than the product of 3 and x is equal to x increased by 15” and the values x = 8, 11, 15, showing work for all three values, tell which is a solution and which is not a solution to the equation. @OrangeMaster

OpenStudy (yttrium):

7 - 3x = x +15

OpenStudy (yttrium):

Just substitute x = 8 then x = 11 ten x = 15 check whether the equality will satisfy then you can conclude that it is a solution, if not, then is is not a solution for example, right we have 7-3x = x + 15 if we let x = 8 7 - 3(8) = 8 + 15 7 - 24 = 23 -17 = 23 --> this not satisfy the equality, therefore, it is not a solution try for the other two!

OpenStudy (anonymous):

I think that I might have messed up and it's: 3x-7=x+15

OpenStudy (anonymous):

Using 8: (3*8)-7=8+15 24-7=23 17=23......Not correct --------------- Using 11: (3*11)-7=11+15 33-7=26 26=26.......CORRECT -------------------- Using 15: (3*15)-7=15+15 45-7=30 38=30......Not correct

OpenStudy (anonymous):

Sorry about getting the first part wrong. It was a tip-off when it didn't work for any of the values. But, my last two comments are correct.